Two minor league baseball players got a total of 222 hits. Washington had 14 more hits than Sanchez. Find the number of hits for each player.
1
Expert's answer
2012-02-16T08:53:56-0500
Let’s denote number of hits of player Washingtonby x. Then the second player had y=x-14hits
Due to the question statement togetherplayers had 222 hits so x+y=222
So we have two equations for x and y:
x+y=222
y=x-14
Substituting expresson for y (secondequation) into the first equation we get
X+y=x+x-14=222
Hence 2x=222+14=236
X=118 – number of Washington’s hits
Number of hits for Sanchez y=x-4=118-4=114
Answer: 118, 114